# Junhan Kim    Ho-young Kim    Eulrang Cho    Chungman Lee    Joonyoung Kim    Yongkweon Jeon

### v2

@@ Abstract
Quantization is a promising solution for deploying large language models (LLMs) on resource-constrained devices. Early methods developed for smaller networks like ResNet rely on gradient-based optimization, which becomes impractical for hyper-scale LLMs with billions of parameters. While recently proposed backpropagation-free or transformation-based methods alleviate this issue, their performance remains limited by either a lack of inter-layer dependency consideration or the use of naïve nearest-rounding-based integer weight assignment to save the heavy computational cost of weight optimization. We thus introduce a novel backpropagation-free PTQ algorithm that optimizes integer weights by considering inter-layer dependencies. The key innovation is the development of attention-aware Hessian matrices that capture inter-layer interactions within the attention module. Extensive experiments demonstrate that our approach not only outperforms existing weight quantization methods but also shows good synergy with conventional methods to suppress activation outliers, leading to state-of-the-art weight-activation quantization performance.
@@ 1 Introduction
Two main categories of quantization approaches have been proposed to preserve the performance of original models: quantization-aware training (QAT) and post-training quantization (PTQ). Although QAT can potentially outperform PTQ, its practicality diminishes when handling hyper-scale LLMs featuring billions of parameters. Consequently, recent quantization efforts have been directed toward PTQ.
Although classic PTQ methods have successfully quantized small models such as ResNet (Nagel et al., 2020; Li et al., 2021), they rely on time-consuming gradient-based optimization, so their efficacy decreases when the complexity of LLMs increases. Accordingly, recent efforts have shifted to develop (a) backpropagation-free methods that optimize integer weights based on Hessian (Frantar et al., 2023) or (b) transformation-based methods that convert a model to be robust to quantization by applying smoothing, rotation, and permutation (Shao et al., 2023; Ashkboos et al., 2024; Liu et al., 2024; Lin et al., 2024a). However, their weight quantization performance is limited as they ignore inter-layer dependencies or use naïve nearest rounding for weight quantization.
In this paper, we propose a novel backpropagation-free PTQ algorithm that considers inter-layer dependencies in optimizing integer weights. Our contributions are as follows:
  • We propose a novel PTQ algorithm called BoA111 Backpropagation-free optimization for Attention-aware PTQ. The key contribution is to exploit the attention reconstruction error, not the layer-wise reconstruction error, in approximating the Hessian to consider inter-layer dependencies within the attention module (Section 3.2). To our knowledge, BoA is the first method to optimize integer weights by considering inter-layer dependencies without relying on gradient-based optimization.
@@ 2 Related Works
Early works aimed to optimize the weight-rounding mechanism (Nagel et al., 2020; Hubara et al., 2021; Li et al., 2021). Instead of allocating the nearest integer, these studies attempted to assign integer weights that minimize the reconstruction error. One popular approach is AdaRound, which learns integer weights satisfying 1 via backpropagation (Nagel et al., 2020). This algorithm has been extended to Brecq where the block-wise reconstruction error has been used, instead of the layer-wise reconstruction error, to consider the inter-layer dependencies (Li et al., 2021). Although AdaRound and Brecq have successfully quantized small-sized models, they rely on time-consuming gradient-based optimizations, which renders their application to LLMs with billions of parameters challenging. Consequently, recent efforts have shifted towards the development of cost-effective quantization methods for LLMs.
@@
These efforts can be classified into two orthogonal classes: (a) backpropagation-free Hessian-based integer weight optimization methods (e.g., GPTQ (Frantar et al., 2023)) and (b) transformation-based methods that convert a model to be more robust to quantization by applying smoothing (e.g., SmoothQuant (Xiao et al., 2023), OmniQuant (Shao et al., 2023), and AffineQuant (Ma et al., 2024)), rotation (e.g., QuaRot (Ashkboos et al., 2024) and SpinQuant (Liu et al., 2024)), or permutation (e.g., DuQuant (Lin et al., 2024a)). The proposed BoA belongs to the class (a) in that it optimizes integer weights using the Hessian without relying on backpropagation. Furthermore, similar to GPTQ, BoA can be combined with transformation-based methods such as QuaRot and SpinQuant (see Section 4). The primary difference over GPTQ lies in our optimization objective: while GPTQ assumes layer-wise independence and focuses on layer-wise reconstruction (which leads to performance degradation), our approach explicitly aims to preserve the attention output, enabling us to account for inter-layer dependencies within the attention module.
@@
Other algorithms exploiting different strategies have also been proposed. For example, SpQR (Dettmers et al., 2023), SqueezeLLM (Kim et al., 2023), and OAC (Edalati et al., 2024) proposed a mixed-precision approach that assigns a large bit-width to quantization-sensitive weights or retains them in full-precision. Compared to the standard uniform quantization, these algorithms require additional processing and memory costs in the inference and need special hardware and dedicated kernels without which accelerating the inference may not be easy. Furthermore, unlike server-grade GPUs, on-device NPUs (e.g. Qualcomm Hexagon) lack support for the mixed precision format, and customizing kernels for desired functionalities is very challenging. Thus, we exclude these algorithms in our comparison and focus on the more universally supported uniform quantization format.
@@ 3.1 Overview of Proposed BoA
$\bm{\delta}{\bm{w}}=\frac{{\mathcal{Q}}(w_{q})-w_{q}}{[\mathbf{U}]_{q,q}}[\mathbf{U}]_{q,:}\text{ where }\mathbf{U}$ $=\operatornamewithlimits{Chol}(\mathbf{H}^{-1})^{T},$ (2)
@@
$\mathbf{H}^{({\bm{w}}^{(\ell)})}$ $\approx 2\mathbf{X}^{(\ell-1)}\mathbf{X}^{(\ell-1)^{T}}\otimes\mathbf{I},$ (3)
where $\mathbf{H}^{({\bm{w}}^{(\ell)})}$ is the Hessian for the $\ell$-th layer, $\otimes$ denotes the Kronecker product operation and $\mathbf{I}$ is the identity matrix. The approximated Hessian $\mathbf{H}^{({\bm{w}}^{(\ell)})}$ in GPTQ relies solely on the input, which means that GPTQ cannot consider the influence of other layers. In other words, GPTQ neglects inter-layer dependencies within the attention module, a crucial aspect of Transformers, which results in limited performance (Jeon et al., 2023). To overcome this, we develop Hessians that incorporate inter-layer dependencies and then use them instead of the conventional Hessian in 3.
@@
0:  weights $\mathbf{W}\in\mathbb{R}^{d_{\text{row}}\times d_{\text{col}}}$ and inputs $\mathbf{X}$ of the Transformer layer
1:  Initialize quantized output: $\mathbf{Q}\leftarrow\mathbf{0}_{H\times d_{\text{row}}/H\times d_{\text{col}}}$
2:  Initialize (row-wise) quantization errors: $\mathbf{E}\leftarrow\mathbf{0}_{H\times d_{\text{col}}}$
3:  Compute attention-aware Hessians: $\mathbf{H}_{h}=\mathbf{H}_{\text{col},h}\otimes\mathbf{H}_{\text{row},h}$                           $\triangleright$ see Table 1
4:  Set step size (scale) $\mathbf{S}$: $\min_{\mathbf{S}}\hskip 0.56905pt\text{tr}\left(\Delta\mathbf{W}\mathbf{H}_{\text{col},h}\Delta\mathbf{W}^{T}\right)$
5:  Compute inverse Hessians $\mathbf{H}_{\text{col},h}^{-1}$ and $\mathbf{H}_{\text{row},h}^{-1}$
6:  Compute $\mathbf{U}_{\text{col},h}=\operatornamewithlimits{Chol}(\mathbf{H}_{\text{col},h}^{-1})^{T}$ and $\mathbf{U}_{\text{row},h}=\operatornamewithlimits{Chol}(\mathbf{H}_{\text{row},h}^{-1})^{T}$
7:  for $j=0,\ldots,d_{\text{row}}/H-1$ do
8:     Construct $\mathbf{W}^{(j)}\in\mathbb{R}^{H\times d_{\text{col}}}$ by stacking the $j$-th rows $[\mathbf{W}_{h}]_{j,:}$
9:     Quantize $\mathbf{W}^{(j)}$: $(\mathbf{Q}_{:,j,:},\mathbf{E})\leftarrow\text{GPTQ}(\mathbf{W}^{(j)},\mathbf{U}_{\text{col},h},\mathbf{S})$                            $\triangleright$ see Appendix D
10:     Update remaining rows: $[\mathbf{W}_{h}]_{j:,:}\leftarrow[\mathbf{W}_{h}]_{j:,:}-\frac{[\mathbf{U}_{\text{row},h}^{T}]_{j:,j}\cdot\mathbf{E}_{h,:}\cdot\mathbf{U}_{\text{col},h}}{[\mathbf{U}_{\text{row},h}]_{j,j}}$    $\triangleright$ see Proposition 3.1
11:  end for
11:  quantized weights $\mathbf{Q}$
@@ 3.2 Proposed Attention-aware Hessian
$\operatornamewithlimits{MHA}(\mathbf{X})=\sum_{h=1}^{H}\mathbf{W}_{\text{out},h}(\mathbf{A}_{h}\mathbf{V}_{h})^{T},~{}\mathbf{A}_{h}=\sigma\left(\frac{\mathbf{Q}_{h}\mathbf{K}_{h}^{T}}{\sqrt{d_{h}}}\right),$
@@
$=\left\|\frac{\mathbf{W}_{\text{out},h}\mathbf{V}_{h}^{T}\mathbf{J}_{\sigma}\mathbf{K}_{h}}{\sqrt{d_{h}}}\Delta\mathbf{W}_{Q,h}\mathbf{X}\right\|_{F}^{2},$ (5)
@@
As in the quantization of the query projection weights, the attention weight $\mathbf{A}_{h}$ changes when quantizing the key projection weights $\mathbf{W}_{K,h}$. By following the steps for 4, $\mathbf{A}_{h}$ can be approximated as
@@
$\|\Delta\operatornamewithlimits{MHA}(\mathbf{X})\|_{F}^{2}=\left\|\frac{\mathbf{Q}_{h}}{\sqrt{d_{h}}}\Delta\mathbf{W}_{K,h}\mathbf{X}\mathbf{J}_{\sigma}^{T}\mathbf{V}_{h}\mathbf{W}_{\text{out},h}^{T}\right\|_{F}^{2}.$
@@ 3.3 Efficient Implementation of BoA
While inter-layer dependencies within the attention module can be considered by exploiting the proposed Hessians, they are significantly more complex than the conventional Hessian in 3, which may incur high computational costs. For example, computing the proposed Hessians in 6 and 7 would be more expensive than computing the conventional one in 3. In this subsection, we present techniques to mitigate the computational overheads incurred by the proposed attention-aware Hessians.
@@
The largest overhead related to the computation of the proposed Hessians is the Jacobian matrix $\mathbf{J}_{\sigma}$ in 6 and 7. For an input sequence of length $L$, the shape of $\mathbf{J}_{\sigma}$ is $H\times L\times L\times L$, which requires a large amount of memory and high computational cost (e.g., more than 400 GB even for the OPT-125M model when $L=2048$).
To mitigate such overhead, we establish a relaxed Hessian that does not require computing $\mathbf{J}_{\sigma}$. To this end, we build the following upper bound for the attention reconstruction error in 5, which will be used as its surrogate:
@@
For the efficient inverse computation, we exploit the useful properties of the Kronecker product (see 13a-13c in Appendix A). Specifically, let $\mathbf{H}=\mathbf{H}_{\text{col}}\otimes\mathbf{H}_{\text{row}}$ where $\mathbf{H}_{\text{col}}\in\mathbb{R}^{d\times d}$ and $\mathbf{H}_{\text{row}}\in\mathbb{R}^{d_{h}\times d_{h}}$, then we obtain
@@
Because the proposed Hessians model the dependency between different rows (e.g., $\mathbf{H}_{\text{row}}=\mathbf{K}_{h}^{T}\mathbf{K}_{h}$ for $\mathbf{W}_{Q,h}$; see 10), we can compensate for the quantization error of a certain row by updating other rows. Specifically, the row-update formula is formulated as in the following proposition for given Hessian $\mathbf{H}=\mathbf{H}_{\text{col}}\otimes\mathbf{H}_{\text{row}}$.
@@
See Appendix B. ∎
Figure 1: Illustration of BoA for the query projection $\mathbf{W}_{Q}$.
Proposition 3.1 implies that the conventional GPTQ cannot compensate quantization error of certain row by updating other rows because $\mathbf{H}_{\text{row}}=\mathbf{I}$ (see 3) and thus $\mathbf{U}_{\text{row}}=\mathbf{I}$ and $[\delta\mathbf{W}]_{i,:}=\mathbf{0}$ for all $i\neq j$.
While the proposed BoA can compensate for the quantization error of each row, the rows must be quantized sequentially (not simultaneously). For example, the second row can be quantized after being updated to compensate for the quantization error of the first row. To accelerate the quantization process, we assume independence between different attention heads (see Figure 1(a)), under which rows related to different heads are independent and can thus be quantized together. For a better understanding, we consider the query projection $\mathbf{W}_{Q}$ as an example (see Figure 1(b)). In the quantization step, we stack the $j$-th rows $[\mathbf{W}_{Q,h}]_{j,:}$ of all different heads, constructing the sub-weight matrix $\mathbf{W}_{Q}^{(j)}\in\mathbb{R}^{H\times d}$ (line 8 in Algorithm 1). Because the rows of $\mathbf{W}_{Q}^{(j)}$ are mutually independent, all the rows of $\mathbf{W}_{Q}^{(j)}$ can be quantized simultaneously as in GPTQ (line 9 in Algorithm 1). Following the quantization of $j$-th rows, we compensate for the quantization error by updating the remaining rows. In this update step, we use the refined weight-update formula in 12 (line 10 in Algorithm 1).
We measure quantization processing times of BoA with and without simultaneous quantization to evaluate how much the head-wise simultaneous quantization can accelerate the quantization process. From Table 2, we observe that BoA requires a significantly long processing time without the simultaneous quantization (more than one day even for the 7B model). This is because all rows need to be quantized sequentially (e.g., 4096 rows are quantized sequentially for LLaMA-7b) and thus the massive compute capabilities of modern GPUs cannot be utilized properly. As evident, we can significantly reduce the processing time by applying the head-wise simultaneous quantization (more than 40 times reduction on the 30B model).
@@
Simultaneous Quantization 7B 13B 30B

### v3

@@ Abstract
Post-training quantization (PTQ) is a promising solution for deploying large language models (LLMs) on resource-constrained devices. Early methods developed for small-scale networks, such as ResNet, rely on gradient-based optimization, which becomes impractical for hyper-scale LLMs with billions of parameters. While recently proposed backpropagation-free or transformation-based methods alleviate this issue, they ignore inter-layer interactions or use the naïve nearest-rounding-based quantized weight assignment to save the heavy computational cost of weight optimization. In this paper, we introduce a novel backpropagation-free PTQ algorithm that optimizes quantized weights by considering inter-layer dependencies. The key innovation is the development of attention-aware Hessian matrices that capture inter-layer interactions within the attention module. Extensive experiments demonstrate that our approach not only outperforms existing weight quantization methods but also shows good synergy with conventional methods to suppress activation outliers, leading to state-of-the-art weight-activation quantization performance. The code will be available at https://github.com/SamsungLabs/BoA.
@@ 1 Introduction
Two main categories of quantization approaches have been proposed to preserve the performance of original models: quantization-aware training (QAT) (Jin et al., 2021; Liu et al., 2023) and post-training quantization (PTQ) (Nagel et al., 2020; Kim et al., 2024). Although QAT can potentially outperform PTQ, its practicality diminishes when handling hyper-scale LLMs featuring billions of parameters. Consequently, recent quantization efforts have been directed toward PTQ.
Although classic PTQ methods have successfully quantized small models such as ResNet (Nagel et al., 2020; Li et al., 2021; Jeon et al., 2022), they rely on time-consuming gradient-based optimization, so their efficacy decreases when the complexity of LLMs increases. Accordingly, recent efforts have shifted to develop (a) backpropagation-free methods that optimize quantized weights based on Hessian (Frantar et al., 2023) or (b) transformation-based methods that convert a model to be robust to quantization by applying smoothing, rotation, and permutation (Shao et al., 2023; Ashkboos et al., 2024; Liu et al., 2024; Lin et al., 2024). However, their weight quantization performance is limited as they ignore inter-layer dependencies or use naïve nearest rounding for weight quantization.
In this paper, we propose a novel backpropagation-free PTQ algorithm that considers inter-layer dependencies in optimizing quantized weights. Our contributions are as follows:
  • We propose a novel PTQ algorithm called BoA111 Backpropagation-free optimization for Attention-aware PTQ. The key contribution is to exploit the attention reconstruction error, not the layer-wise reconstruction error, in approximating the Hessian to consider inter-layer dependencies within the attention module (Section 3.2). To our knowledge, BoA is the first method to optimize quantized weights by considering inter-layer dependencies without relying on gradient-based optimization.
@@ 2 Related Works
Early works aimed to optimize the weight-rounding mechanism (Nagel et al., 2020; Hubara et al., 2021; Li et al., 2021; Jeon et al., 2023a). Instead of allocating the nearest integer, these studies attempted to assign integer weights that minimize the reconstruction error. One popular approach is AdaRound, which learns quantized weights satisfying 1 via backpropagation (Nagel et al., 2020). This algorithm has been extended to Brecq where the block-wise reconstruction error has been used, instead of the layer-wise reconstruction error, to consider the inter-layer dependencies (Li et al., 2021). Although AdaRound and Brecq have successfully quantized small-sized models, they rely on time-consuming gradient-based optimizations, which renders their application to LLMs with billions of parameters challenging. Consequently, recent efforts have shifted towards the development of cost-effective quantization methods for LLMs.
@@
These efforts can be classified into two orthogonal classes: (a) backpropagation-free Hessian-based integer weight optimization methods (e.g., GPTQ (Frantar et al., 2023)) and (b) transformation-based methods that convert a model to be more robust to quantization by applying smoothing (e.g., SmoothQuant (Xiao et al., 2023), OmniQuant (Shao et al., 2023), and AffineQuant (Ma et al., 2024)), rotation (e.g., QuaRot (Ashkboos et al., 2024) and SpinQuant (Liu et al., 2024)), or permutation (e.g., DuQuant (Lin et al., 2024)).
The proposed BoA belongs to the class (a) in that it optimizes quantized weights using the Hessian without relying on backpropagation. Furthermore, similar to GPTQ, BoA can be combined with transformation-based methods such as QuaRot and SpinQuant (see Section 4). The primary difference over GPTQ lies in our optimization objective: while GPTQ assumes layer-wise independence and focuses on layer-wise reconstruction (which leads to performance degradation), our approach explicitly aims to preserve the attention output, enabling us to account for inter-layer dependencies within the attention module.
@@
Other algorithms exploiting different strategies have also been proposed. For example, SpQR (Dettmers et al., 2023), SqueezeLLM (Kim et al., 2023), and OAC (Edalati et al., 2024) proposed a mixed-precision approach that assigns a large bit-width to quantization-sensitive weights or retains them in full-precision. Compared to the standard uniform quantization, these algorithms require additional processing and memory costs in the inference and need special hardware and dedicated kernels without which accelerating the inference may not be easy. Furthermore, unlike server-grade GPUs, on-device NPUs (e.g. Qualcomm Hexagon) lack support for the mixed precision format, and customizing kernels for desired functionalities is very challenging. Thus, we exclude these algorithms in our comparison and focus on the more universally supported uniform quantization format. We also exclude recent vector quantization approaches (e.g. QuIP# (Tseng et al., 2024) and AQLM (Egiazarian et al., 2024)) because they need additional memory (bits) for storing codebooks, which are required to perform the dequantization during the inference.
@@ 3.1 Overview of Proposed BoA
$\bm{\delta}\mathbf{w}=\frac{{\mathcal{Q}}(w_{q})-w_{q}}{[\mathbf{U}]_{q,q}}[\mathbf{U}]_{q,:}\text{ where }\mathbf{U}$ $=\operatornamewithlimits{Chol}(\mathbf{H}^{-1})^{T},$ (2)
@@
$\mathbf{H}^{(\mathbf{w}^{(\ell)})}$ $\approx 2\mathbf{X}^{(\ell-1)}\mathbf{X}^{(\ell-1)^{T}}\otimes\mathbf{I},$ (3)
where $\mathbf{w}^{(\ell)}$ is the flattened representation of $\mathbf{W}^{(\ell)}$, $\mathbf{H}^{(\mathbf{w}^{(\ell)})}$ is the Hessian for the $\ell$-th layer, $\otimes$ denotes the Kronecker product operation, and $\mathbf{I}$ is the identity matrix. The approximated Hessian $\mathbf{H}^{(\mathbf{w}^{(\ell)})}$ in GPTQ relies solely on the input, which means that GPTQ cannot consider the influence of other layers. In other words, GPTQ neglects inter-layer dependencies within the attention module, a crucial aspect of Transformers, which results in limited performance (Jeon et al., 2023b). To overcome this, we develop Hessians that incorporate inter-layer dependencies and then use them instead of the conventional Hessian in 3.
@@
    For $\mathbf{W}_{Q,h}$ and $\mathbf{W}_{K,h}$ of models exploiting rotary position embedding, see 14 and 15.
@@
0:  weights $\mathbf{W}_{\{Q,K,V\}}\in\mathbb{R}^{H\times d_{h}\times d}$ and inputs $\mathbf{X}$ of the Transformer layer
1:  for $\mathbf{W}\in\{\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}\}$ do
2:     Initialize quantized output: $\mathbf{Q}\leftarrow\mathbf{0}_{H\times d_{h}\times d}$
3:     Initialize (row-wise) quantization errors: $\mathbf{E}\leftarrow\mathbf{0}_{H\times d}$
4:     Compute attention-aware Hessians: $\mathbf{H}_{h}=\mathbf{H}_{\text{col},h}\otimes\mathbf{H}_{\text{row},h}$                           $\triangleright$ see Table 1
5:     Set step size (scale) $\mathbf{S}$: $\min_{\mathbf{S}}\hskip 0.56905pt\text{tr}\left(\Delta\mathbf{W}_{h}\mathbf{H}_{\text{col},h}\Delta\mathbf{W}_{h}^{T}\right)$
6:     Compute inverse Hessians $\mathbf{H}_{\text{col},h}^{-1}$ and $\mathbf{H}_{\text{row},h}^{-1}$
7:     Compute $\mathbf{U}_{\text{col},h}=\operatornamewithlimits{Chol}(\mathbf{H}_{\text{col},h}^{-1})^{T}$ and $\mathbf{U}_{\text{row},h}=\operatornamewithlimits{Chol}(\mathbf{H}_{\text{row},h}^{-1})^{T}$
8:     for $j=0,\ldots,d_{h}-1$ do
9:        Construct $\mathbf{W}^{(j)}\in\mathbb{R}^{H\times d}$ by stacking the $j$-th rows $[\mathbf{W}_{h}]_{j,:}$
10:        Quantize $\mathbf{W}^{(j)}$: $(\mathbf{Q}_{:,j,:},\mathbf{E})\leftarrow\text{GPTQ}(\mathbf{W}^{(j)},\mathbf{U}_{\text{col},h},\mathbf{S})$                            $\triangleright$ see Appendix E
11:        Update remaining rows: $[\mathbf{W}_{h}]_{j:,:}\leftarrow[\mathbf{W}_{h}]_{j:,:}-\frac{[\mathbf{U}_{\text{row},h}^{T}]_{j:,j}\cdot\mathbf{E}_{h,:}\cdot\mathbf{U}_{\text{col},h}}{[\mathbf{U}_{\text{row},h}]_{j,j}}$     $\triangleright$ see Proposition 3.1
12:     end for
13:  end for
13:  quantized weights $\mathbf{Q}$
@@ 3.2 Proposed Attention-aware Hessian
$\operatornamewithlimits{MHA}(\mathbf{X})=\sum_{h=1}^{H}\mathbf{W}_{\text{out},h}(\mathbf{A}_{h}\mathbf{V}_{h})^{T},\mathbf{A}_{h}=\sigma\left(\frac{\mathbf{Q}_{h}\mathbf{K}_{h}^{T}}{\sqrt{d_{h}}}\right),$ (4)
@@
$\approx\left\|\frac{\mathbf{W}_{\text{out},h}\mathbf{V}_{h}^{T}\mathbf{J}_{\sigma}\mathbf{K}_{h}}{\sqrt{d_{h}}}\Delta\mathbf{W}_{Q,h}\mathbf{X}\right\|_{F}^{2},$ (6)
@@
As in the quantization of the query projection weights, the attention weight $\mathbf{A}_{h}$ changes when quantizing the key projection weights $\mathbf{W}_{K,h}$. By following the steps for 5, $\mathbf{A}_{h}$ can be approximated as
@@
$\|\Delta\operatornamewithlimits{MHA}(\mathbf{X})\|_{F}^{2}\approx\left\|\frac{\mathbf{Q}_{h}}{\sqrt{d_{h}}}\Delta\mathbf{W}_{K,h}\mathbf{X}\mathbf{J}_{\sigma}^{T}\mathbf{V}_{h}\mathbf{W}_{\text{out},h}^{T}\right\|_{F}^{2}.$
@@ 3.3 Efficient Implementation of BoA
While inter-layer dependencies within the attention module can be considered by exploiting the proposed Hessians, they are significantly more complex than the conventional Hessian in 3, which may incur high computational costs. For example, computing the proposed Hessians in 7 and 8 would be more expensive than computing the conventional one in 3. In this subsection, we present techniques to mitigate the computational overheads incurred by the proposed attention-aware Hessians.
@@
The largest overhead related to the computation of the proposed Hessians is the Jacobian matrix $\mathbf{J}_{\sigma}$ in 7 and 8. For an input sequence of length $L$, the shape of $\mathbf{J}_{\sigma}$ is $H\times L\times L\times L$, which requires a large amount of memory and high computational cost (e.g., more than 400 GB even for the OPT-125M model when $L=2048$).
To mitigate such overhead, we establish a relaxed Hessian that does not require computing $\mathbf{J}_{\sigma}$. To this end, we build the following upper bound for the attention reconstruction error in 6, which will be used as its surrogate:
@@
We note that if rotary position embedding (RoPE) (Su et al., 2023) is used, the MHA output is different from that in 4, and thus the corresponding attention-aware Hessians would also be different. Specifically, since the attention weight $\mathbf{A}_{h}$ for models exploiting RoPE is expressed as
$\mathbf{A}_{h}=\sigma\left(\frac{\widetilde{\mathbf{Q}}_{h}\widetilde{\mathbf{K}}_{h}^{T}}{\sqrt{d_{h}}}\right),\widetilde{\mathbf{Q}}_{h}=\operatornamewithlimits{RoPE}(\mathbf{Q}_{h}),\widetilde{\mathbf{K}}_{h}=\operatornamewithlimits{RoPE}(\mathbf{K}_{h}),$ (13)
the surrogate $\|\mathbf{K}_{h}\Delta\mathbf{Q}_{h}^{T}\|_{F}^{2}$ used to develop the attention-aware Hessian in 11 changes to $\|\widetilde{\mathbf{K}}_{h}\Delta\widetilde{\mathbf{Q}}_{h}^{T}\|_{F}^{2}$. As a result, we obtain different Hessians for models exploiting RoPE (see Appendix B for the detailed derivation):
$\mathbf{H}^{(\mathbf{w}_{Q,h})}$ $=2\mathbf{X}\mathbf{X}^{T}\otimes\frac{1}{L}\sum_{\ell=1}^{L}\mathbf{R}_{\ell}^{T}\widetilde{\mathbf{K}}_{h}^{T}\widetilde{\mathbf{K}}_{h}\mathbf{R}_{\ell},$ (14)
$\mathbf{H}^{(\mathbf{w}_{K,h})}$ $=2\mathbf{X}\mathbf{X}^{T}\otimes\frac{1}{L}\sum_{\ell=1}^{L}\mathbf{R}_{\ell}^{T}\widetilde{\mathbf{Q}}_{h}^{T}\widetilde{\mathbf{Q}}_{h}\mathbf{R}_{\ell},$ (15)
where $\mathbf{R}_{\ell}$ be the rotary matrix for the $\ell$-th token (see eq. (15) in (Su et al., 2023)).
@@
For the efficient inverse computation, we exploit the useful properties of the Kronecker product (see 17a-17c in Appendix A). Specifically, let $\mathbf{H}=\mathbf{H}_{\text{col}}\otimes\mathbf{H}_{\text{row}}$ where $\mathbf{H}_{\text{col}}\in\mathbb{R}^{d\times d}$ and $\mathbf{H}_{\text{row}}\in\mathbb{R}^{d_{h}\times d_{h}}$, then we obtain
@@
Because the proposed Hessians model the dependency between different rows (e.g., $\mathbf{H}_{\text{row}}=\mathbf{K}_{h}^{T}\mathbf{K}_{h}$ for $\mathbf{W}_{Q,h}$; see 11), we can compensate for the quantization error of a certain row by updating other rows. Specifically, the row-update formula is formulated as in the following proposition for given Hessian $\mathbf{H}=\mathbf{H}_{\text{col}}\otimes\mathbf{H}_{\text{row}}$.
@@
See Appendix C. ∎
The update formula in 16 implies that we do not need to compute and store the Cholesky decomposition $\mathbf{U}_{h}=\mathbf{U}_{\text{col},h}\otimes\mathbf{U}_{\text{row},h}$ of the full Hessian $\mathbf{H}_{h}$ for updating weights; only $\mathbf{U}_{\text{col},h}$ and $\mathbf{U}_{\text{row},h}$ are sufficient for updating weights, and thus we can save the memory cost caused by the high-dimensional Kronecker product operation. Proposition 3.1 also implies that the conventional GPTQ cannot compensate quantization error of certain row by updating other rows because $\mathbf{H}_{\text{row}}=\mathbf{I}$ (see 3) and thus $\mathbf{U}_{\text{row}}=\mathbf{I}$ and $[\delta\mathbf{W}]_{i,:}=\mathbf{0}$ for all $i\neq j$.
While the proposed BoA can compensate for the quantization error of each row, the rows must be quantized sequentially (not simultaneously). For example, the second row can be quantized after being updated to compensate for the quantization error of the first row. To accelerate the quantization process, we assume independence between different attention heads (see Figure 1(a)), under which rows related to different heads are independent and can thus be quantized together. For a better understanding, we consider the query projection $\mathbf{W}_{Q}$ as an example (see Figure 1(b)). In the quantization step, we stack the $j$-th rows $[\mathbf{W}_{Q,h}]_{j,:}$ of all different heads, constructing the sub-weight matrix $\mathbf{W}_{Q}^{(j)}\in\mathbb{R}^{H\times d}$ (line 8 in Algorithm 1). Because the rows of $\mathbf{W}_{Q}^{(j)}$ are mutually independent, all the rows of $\mathbf{W}_{Q}^{(j)}$ can be quantized simultaneously as in GPTQ (line 9 in Algorithm 1). Following the quantization of $j$-th rows, we compensate for the quantization error by updating the remaining rows. In this update step, we use the refined weight-update formula in 16 (line 10 in Algorithm 1).
@@
Simultaneous Quantization LLaMA Model Size
7B 13B 30B